OpenStudy (anonymous):

a triangular sail has a perimeter of 25 m. side a is 2 m shorter than twice side b, and side c is 3 m longer than side b. find the length of each side.

OpenStudy (anonymous):

P=25 a+b+c=25 a=2b-2, c=b+3 (2b-2) + (b) + (b+3) = 25 2b -2 + b + b + 3 = 25 4b = 24 b = 6 a = 2b-2 c = b + 3 a = 2(6) - 2 c = 6 + 3 a = 10 c = 9 Check: a + b + c = 25 10 + 6 + 9 ?= 25, yes this is true